# Disk patterns, quasi-duality and the uniform bounded diameter conjecture

### v1

@@ 2.5. Acylindrical reflection groups
Given $(\mathcal{G},\omega)$, let $F$ be a hyperbolic face, and $v,w$ two nonadjacent vertices on $\partial F$. If there exists a vertex $x\notin\partial F$ so that $\omega(xv)=\omega(xw)=\pi/2$, we call the path $vxw$ a right-angled 2-connection. We aim to prove the following statement relating connectedness of $(\mathcal{G},\omega)$ with acylindricity of the group $\widetilde{G}$. This generalizes Proposition 3.6 of [LLM22].
@@
For the other direction, we construct a new plane graph $(\widehat{\mathcal{G}},\widehat{\omega})$ as follows. For each hyperbolic face $F$ of $\mathcal{G}$, we add a new vertex $v_{F}$ and connect it to all vertices on $\partial F$. Moreover, we define the weight $\widehat{\omega}$ on these new edges to be $\pi/2$, and the same as $\omega$ otherwise.
Note that the new graph $(\widehat{\mathcal{G}},\widehat{\omega})$ has no hyperbolic face. Indeed, any face $F$ of $\widehat{\mathcal{G}}$ is either a face of $\mathcal{G}$ (in which case it is already elliptic or parabolic), or a triangle formed by $v_{F^{\prime}}$ and two adjacent vertices $v,w$ of some face $F^{\prime}$ of $\mathcal{G}$. Since $\widehat{\omega}(vv_{F^{\prime}})+\widehat{\omega}(wv_{F^{\prime}})+\widehat{\omega}(vw)\geq\pi$, the face $F$ is elliptic or parabolic.
@@
The group $\widetilde{G}$ is acylindrical if and only if $(\widehat{\mathcal{G}},\widehat{\omega})$ is realizable.
@@
This means that for any hyperbolic face $F$ of $(\mathcal{G},\omega)$, we can add a circle perpendicular to all $D_{v}$ with $v\in\partial F$ – this circle is the boundary of the corresponding component $\Omega_{F}$ of the domain of discontinuity. This is exactly equivalent to $(\widehat{\mathcal{G}},\widehat{\omega})$ being realizable, as desired. ∎
@@
If $\mathcal{G}$ contains exactly four vertices, then it must be a tetrahedron. Suppose further that $(\mathcal{G},\omega)$ has only one hyperbolic face. Then $\widehat{\mathcal{G}}$ contains exactly 5 vertices. Since $\widetilde{G}$ is acylindrical if and only if $(\widehat{\mathcal{G}},\widehat{\omega})$ is realizable, we can apply apply Proposition 2.4 to conclude that the condition in Part (2) is necessary and sufficient.
For the remainder of the proof, we assume $(\mathcal{G},\omega)$ contains more than four vertices or has at least 2 hyperbolic faces. Then $(\widehat{\mathcal{G}},\widehat{\omega})$ has at least 6 vertices.
First suppose $\widetilde{G}$ is acylindrical. Then $(\widehat{\mathcal{G}},\widehat{\omega})$ is realizable. We claim that $(\mathcal{G},\omega)$ cannot contain any right-angled 2-connection. Suppose otherwise, and let $vxw$ be a right-angled 2-connection with $v,w\in\partial F$ for some hyperbolic face $F$. Then $v,x,w,v_{F}$ form a 4-cycle in $\widehat{\mathcal{G}}$ with weights on edges between them add up to $2\pi$. This is impossible by Theorem 2.3, as $v,w$ are assumed to be nonadjacent, and $v_{F}$ and $x$ are nonadjacent by construction.
Conversely, suppose the 3-connected graph $(\mathcal{G},\omega)$ contains no right-angled 2-connections. We need to check that $(\widehat{\mathcal{G}},\widehat{\omega})$ satisfy all conditions of Theorem 2.3. First note that any new 3-cycles bound a new face, so the first condition is met. Furthermore, any new 4-cycle must be of one of the following two possibilities.
@@
Hence $(\widehat{\mathcal{G}},\widehat{\omega})$ is realizable, and thus $\widetilde{G}$ is acylindrical, as desired. ∎
@@
(a) A 3-connected graph. The weights on all edges except the red ones are 00.
(b) The weights on the red edges are 00; this is acylindrical, and the limit set is homeomorphic to a circle packing.
@@ 3.1. Vertex extremal length / width
[…]
@@ 3.2. Polygonal subdivision graph
Recall that a CW complex $Y$ is a subdivision of a CW complex $X$ if $X=Y$ and every closed cell of $Y$ is contained in a closed cell of $X$. We define a polygon $P$ as a finite CW complex homeomorphic to a closed disk that contains one 2-cell, with at least three 0-cells. $P$ is called $n$-gon if it has $n$ 0-cells. We will also call 0-cells, 1-cells and 2-cells the vertices, edges and faces respectively.
@@ 3.3. Triangulation $\widetilde{\mathcal{G}}$
Denote the vertex set of $\mathcal{G}$ and $\widetilde{\mathcal{G}}$ by $\mathcal{V}$ and $\widetilde{\mathcal{V}}$ respectively. Let $\widetilde{\mu}\colon\widetilde{\mathcal{V}}\longrightarrow[0,\infty)$ be an extremal $\widetilde{\Gamma}_{a,b}$-admissible metric on $\widetilde{\mathcal{G}}$ relative to $\partial P$. We remark that the existence of the extremal metric follows from compactness of admissible metrics. We define $\mu$ to be the restriction of $\widetilde{\mu}$ on $\mathcal{V}\subseteq\widetilde{\mathcal{V}}$. Let $\gamma$ be a proper path in $\mathcal{G}$ that connects $a,b$. Then $\gamma$ is also a proper path in $\widetilde{\mathcal{G}}$ that connects $a,b$, so $l_{\widetilde{\mu}}(\gamma)\geq 1$. Thus, $l_{\mu}(\gamma)=l_{\widetilde{\mu}}(\gamma)\geq 1$. Therefore, $\mu$ is $\Gamma_{a,b}$-admissible metric on $\mathcal{G}$ relative to $\partial P$. Note that $\operatorname{area}(\mu)\leq\operatorname{area}(\widetilde{\mu})$. Hence,
@@ 3.5.1. The setup
Abusing the notations, we use $\mathcal{F}$ to denote the space of non-triangular faces of $\mathcal{R}_{k}(P)$ for any $k=0,...,r$. We denote the additional vertices in $\mathcal{G}_{k}$ by $w_{v,F}$, where $F\in\mathcal{F}$ is adjacent to $v$ and $v\in\{v_{1},...,v_{r}\}$. Note that there is a natural quotient map $q:\mathcal{G}_{r}\longrightarrow\widetilde{\mathcal{G}}$, which collapses the vertices $w_{v,F}$ to $w_{F}$, for $F\in\mathcal{F}$ (see Figure 3.2). It is easy to see that the projection satisfies the following property.
@@ 3.5.2. A sequence of admissible metrics $\mu_{k}$
$\operatorname{area}(\widetilde{\mu})$ $=\sum_{k=1}^{r}\widetilde{\mu}^{2}(v_{k})+\sum_{F\in\mathcal{F}}\widetilde{\mu}^{2}(w_{F})$
@@
Figure 3.3. An illustration of the setup of the construction of the metric $\mu_{k}$.
If $v,w$ are two vertices in $\mathcal{G}_{k-1}$, we use $[v,w]=[v,w]_{k-1}$ to denote a geodesic (with respect to the metric $\mu_{k-1}$) that connects $v,w$. Note that the geodesic $[v,w]$ may not be unique. We will often drop the subscript $k-1$ if the underlying graph $\mathcal{G}_{k-1}$ is not ambiguous. We denote
@@
Consider the set of all vertices $x_{i},i\in S$ with $\mathring{d}_{i}<d$. Then by Lemma 3.10, any geodesic connecting $x_{i}$ to $a$ does not pass through $x$. Consider the set of all geodesics connecting $a$ to $x_{i},i\in S$. Since the graph $\mathcal{G}_{k-1}$ is embedded in the plane, by relabeling the indices if necessary, we denote the right most geodesic not passing through $x$ by $[x_{1},a]$ and the left most one by $[x_{l},a]$. Then $x_{2},...,x_{l-1}$ are all contained in the region bounded by $[a,x_{1}]\cup[x_{1},x]\cup[x,x_{l}]\cup[x_{l},a]$. Note that by construction, we have
@@
Then $\mathring{d}_{j_{0}}\leq\mathring{d}_{j}$ for all $j=1,...,l$. Inductively, for $n\geq 1$, we define
@@
$\mu_{k}(v)=\begin{cases}\mu_{k-1}(v),&\text{ if $v\in\mathcal{V}_{k-1}$}\\ \max\{0,\mathring{d}_{j_{n+1}}-d_{j_{n}}\},&\text{ if $v=y_{j_{n}},n=0,...,q$ and $j_{n}\in\mathcal{I}$}\\ \max\{0,\mathring{d}_{j_{n}}-d_{j_{n+1}}\},&\text{ if $v=y_{j_{n+1}-1},n=p-1,...,-1$ and $j_{n}\in\mathcal{I}$}\\ 0,&\text{ otherwise.}\end{cases}$
@@
Suppose that $\mathring{d}_{i}^{\prime}+e_{i}^{\prime}=d^{\prime}_{i}+\mathring{e}^{\prime}_{i}\geq 1$ for all $i=1,...,s$. Then $\mu_{k}$ is $\Gamma_{a,b,k}$-admissible.
@@
$f(i)\coloneqq\begin{cases}\mathring{d}_{j_{n}},&\text{if $j_{n}\leq i<j_{n+1}$and $n=p,...,-1$}\\ \mathring{d}_{j_{0}},&\text{if $i=j_{0}$}\\ \mathring{d}_{j_{n+1}},&\text{if $j_{n}<i\leq j_{n+1}$ and $n=0,...,q-1$}\\ d,&\text{if $i>j_{q}=l$.}\end{cases}$
@@
If $f(i)=f(j)$, then the inequality holds trivially. Thus, we assume $f(i)>f(j)$. In particular, $f(j)\neq d$, so $j\leq l=j_{q}$. For simplicity of the presentation, we assume that $j_{n}<j\leq j_{n+1}$ for some $n=0,...,q-1$. The other case where $j\leq j_{0}$ can be proved similarly.
Suppose that $\gamma$ is contained in $\mathcal{G}_{k-1}$. Let $[x_{j_{n}},a]_{k-1}$ and $[x_{j_{n+1}},a]_{k-1}$ be geodesics in $\mathcal{G}_{k-1}$ connecting $x_{j_{n}}$ and $x_{j_{n+1}}$ to $a$ respectively. Since $f(i)>f(j)=\mathring{d}_{j_{n+1}}\geq\mathring{d}_{j_{n}}$, $[a,x_{j_{n}}]_{k-1}\cup[x_{j_{n}},x]_{k-1}\cup[x,x_{j_{n+1}}]_{k-1}\cup[x_{j_{n+1}},a]_{k-1}$ separates $x_{j}$ from $x_{i}$. Since $\operatorname{Int}(\gamma)$ contains no vertex in $\{x,x_{1},...,x_{s}\}$, $\gamma$ cuts either $[x_{j_{n}},a]_{k-1}$ or $[x_{j_{n+1}},a]_{k-1}$. Let $\gamma^{\prime}$ be the subpath of $\gamma$ that connects $x_{i}$ to a vertex $v\in[x_{j_{n}},a]_{k-1}\cup[x_{j_{n+1}},a]_{k-1}$. Let $[a,v)_{k-1}$ be the subpath of either $[x_{j_{n}},a]_{k-1}$ or $[x_{j_{n+1}},a]_{k-1}$ connecting $a$ and $v$ but excluding the vertex $v$. Note that $[a,v)_{k-1}\cup\gamma^{\prime}$ is a proper path in $\mathcal{G}_{k-1}$ connecting $a,x_{i}$. Thus by Equation (3.16),
@@
Let $[x_{i},a]_{k}$ be a geodesic in $\mathcal{G}_{k}$ connecting $x_{i}$ and $a$. If $x\in[x_{i},a]_{k}$, then $\mathring{d}_{i}^{\prime}\geq d\geq f(i)$. Therefore, we may assume that $x\notin[x_{i},a]_{k}$. Then we can break $[x_{i},a]_{k}=\gamma_{1}\cup\gamma_{2}\cup...\cup\gamma_{t}$ into finitely many pieces so that $\operatorname{Int}(\gamma_{j})$ contains no vertices in $\{x,x_{1},...,x_{s}\}$ and $\partial\gamma_{j}\in\{a,x_{1},...,x_{s}\}$.
@@
Suppose $i=l+1,...,s$. By Lemma 3.15, we have that $\mathring{d}^{\prime}_{i}\geq d^{\prime}=d$. Since $e^{\prime}_{i}\geq\mathring{e}^{\prime}=\mathring{e}$, and $d+\mathring{e}\geq 1$, we conclude that $\mathring{d}^{\prime}_{i}+e^{\prime}_{i}\geq 1$.
Otherwise, let $[x_{i},b]_{k}$ be the geodesic in $\mathcal{G}_{k}$ connecting $x_{i}$ and $b$. If $x\in[x_{i},b]_{k}$, then $\mathring{e}_{i}^{\prime}\geq e$. Since $d_{i}^{\prime}\geq\mathring{d}$, we conclude that $\mathring{d}^{\prime}_{i}+e^{\prime}_{i}\geq 1$. Therefore, we may assume that $x\notin[x_{i},b]_{k}$. Then we can break $[x_{i},b]_{k}=\gamma_{1}\cup\gamma_{2}\cup...\cup\gamma_{t}$ into finitely many pieces so that $\operatorname{Int}(\gamma_{j})$ contains no vertices in $\{x,x_{1},...,x_{s}\}$ and $\partial\gamma_{j}\in\{b,x_{1},...,x_{s}\}$. We assume that $b\in\partial\gamma_{t}$.
@@
Otherwise, either $j_{n}<i\leq j_{n+1}$ for some $n=0,...,q-1$, or $j_{n}\leq i<j_{n+1}$ for some $n=p,...,-1$. Without loss of generality, we assume that we are in the first case. Let $[x_{j_{n}},a]_{k-1}$ and $[x_{j_{n+1}},a]_{k-1}$ be the geodesics in $\mathcal{G}_{k-1}$ connecting $x_{j_{n}}$ and $x_{j_{n+1}}$ to $a$ respectively. Note that $\mathring{d}_{j_{n}},\mathring{d}_{j_{n+1}}\leq\mathring{d}_{l}<d$ by Equation (3.13). Thus by Lemma 3.10, $x\notin[x_{j_{n}},a]_{k-1}\cup[x_{j_{n+1}},a]_{k-1}$. Then $\gamma$ must intersect either $[x_{j_{n}},a]_{k-1}$ or $[x_{j_{n+1}},a]_{k-1}$. Since $\mu_{k-1}$ is $\Gamma_{a,b,k-1}$-admissible, either $\mathring{d}_{j_{n}}+e_{i}^{\prime}\geq 1$ or $\mathring{d}_{j_{n+1}}+e_{i}^{\prime}\geq 1$. Thus,

### v2

@@ 2.5. Acylindrical reflection groups
Given $(\mathcal{G},\omega)$, let $F$ be a hyperbolic face, and $v,w$ two nonadjacent vertices on $\partial F$. If there exists a vertex $x\notin\partial F$ so that $\omega(xv)=\omega(xw)=\pi/2$, we call the path vxwvxw a right-angled 2-connection. We aim to prove the following statement relating connectedness of $(\mathcal{G},\omega)$ with acylindricity of the group $\widetilde{G}$. This generalizes Proposition 3.6 of [LLM22].
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
(a) A 3-connected graph. The weights on all edges except the red ones are 0.
(b) The weights on the red edges are 0; this is acylindrical, and the limit set is homeomorphic to a circle packing.
@@ 3.1. Vertex extremal length / width
$\operatorname{EW}(\Gamma)=\operatorname{EW}_{\mathcal{G}}(\Gamma)\coloneqq\inf\{\operatorname{area}(\mu)\colon\mu\text{ is $\Gamma$-admissible}\}.$
@@ 3.2. Polygonal subdivision graph
Recall that a CW complex $Y$ is a subdivision of a CW complex $X$ if the underlying topological spaces (which for simplicity we still denote by $X$ and $Y$) are the same, and every closed cell of $Y$ is contained in a closed cell of $X$. We define a polygon $P$ as a finite CW complex homeomorphic to a closed disk that contains one 2-cell, with at least three 0-cells. $P$ is called $n$-gon if it has $n$ 0-cells. We will also call 0-cells, 1-cells and 2-cells the vertices, edges and faces respectively.
@@ 3.3. Triangulation $\widetilde{\mathcal{G}}$
[…]
@@ 3.5.1. The setup
Abusing the notations, we use $\mathcal{F}$ to denote the space of non-triangular faces of $\mathcal{R}_{k}(P)$ for any k=0,,rk=0,...,r. We denote the additional vertices in $\mathcal{G}_{k}$ by $w_{v,F}$, where $F\in\mathcal{F}$ is adjacent to $v$ and $v\in\{v_{1},...,v_{r}\}$. Note that there is a natural quotient map $q:\mathcal{G}_{r}\longrightarrow\widetilde{\mathcal{G}}$, which collapses the vertices $w_{v,F}$ to $w_{F}$, for $F\in\mathcal{F}$ (see Figure 3.2). It is easy to see that the projection satisfies the following property.
@@ 3.5.2. A sequence of admissible metrics $\mu_{k}$
[…]
@@
Figure $3.$ An illustration of the setup of the construction of the metric $\mu_{k}$.
If $v,w$ are two vertices in $\mathcal{G}_{k-1}$, we use $v,w]=[v,w]_{k-1}$ to denote a geodesic (with respect to the metric $\mu_{k-1}$) that connects $v,w$. Note that the geodesic $[v,w]$ may not be unique. We will often drop the subscript k1k-1 if the underlying graph $\mathcal{G}_{k-1}$ is not ambiguous. We denote
@@
Consider the set of all vertices $x_{i},i\in S$ with $\mathring{d}_{i}<d$. Then by Lemma 3.10, any geodesic connecting $x_{i}$ to $a$ does not pass through $x$. Consider the set of all geodesics connecting $a$ to $x_{i},i\in S$. Since the graph $\mathcal{G}_{k-1}$ is embedded in the plane, by relabeling the indices if necessary, we denote the right most geodesic not passing through $x$ by $x_{1},a]$ and the left most one by $x_{l},a]$. Then $x_{2},...,x_{l-1}$ are all contained in the region bounded by $a,x_{1}]\cup[x_{1},x]\cup[x,x_{l}]\cup[x_{l},a]$. Note that by construction, we have
@@
Then $\mathring{d}_{j_{0}}\leq\mathring{d}_{j}$ for all j=1,,lj=1,...,l. Inductively, for $n\geq 1$, we define
@@
[…]
@@
Suppose that $\mathring{d}_{i}^{\prime}+e_{i}^{\prime}=d^{\prime}_{i}+\mathring{e}^{\prime}_{i}\geq 1$ for all i=1,,si=1,...,s. Then $\mu_{k}$ is $\Gamma_{a,b,k}$-admissible.
@@
[…]
@@
If $f(i)=f(j)$, then the inequality holds trivially. Thus, we assume $f(i)>f(j)$. In particular, $f(j)\neq d$, so $j\leq l=j_{q}$. For simplicity of the presentation, we assume that $j_{n}<j\leq j_{n+1}$ for some n=0,,q1n=0,...,q-1. The other case where $j\leq j_{0}$ can be proved similarly.
Suppose that $\gamma$ is contained in $\mathcal{G}_{k-1}$. Let $x_{j_{n}},a]_{k-1}$ and $x_{j_{n+1}},a]_{k-1}$ be geodesics in $\mathcal{G}_{k-1}$ connecting $x_{j_{n}}$ and $x_{j_{n+1}}$ to $a$ respectively. Since $f(i)>f(j)=\mathring{d}_{j_{n+1}}\geq\mathring{d}_{j_{n}}$, $a,x_{j_{n}}]_{k-1}\cup[x_{j_{n}},x]_{k-1}\cup[x,x_{j_{n+1}}]_{k-1}\cup[x_{j_{n+1}},a]_{k-1}$ separates $x_{j}$ from $x_{i}$. Since $\operatorname{Int}(\gamma)$ contains no vertex in $\{x,x_{1},...,x_{s}\}$, $\gamma$ cuts either $x_{j_{n}},a]_{k-1}$ or $x_{j_{n+1}},a]_{k-1}$. Let $\gamma^{\prime}$ be the subpath of $\gamma$ that connects $x_{i}$ to a vertex $v\in[x_{j_{n}},a]_{k-1}\cup[x_{j_{n+1}},a]_{k-1}$. Let $a,v)_{k-1}$ be the subpath of either $x_{j_{n}},a]_{k-1}$ or $x_{j_{n+1}},a]_{k-1}$ connecting $a$ and $v$ but excluding the vertex $v$. Note that $a,v)_{k-1}\cup\gamma^{\prime}$ is a proper path in $\mathcal{G}_{k-1}$ connecting $a,x_{i}$. Thus by Equation (3.16),
@@
Let $x_{i},a]_{k}$ be a geodesic in $\mathcal{G}_{k}$ connecting $x_{i}$ and $a$. If $x\in[x_{i},a]_{k}$, then $\mathring{d}_{i}^{\prime}\geq d\geq f(i)$. Therefore, we may assume that $x\notin[x_{i},a]_{k}$. Then we can break $x_{i},a]_{k}=\gamma_{1}\cup\gamma_{2}\cup...\cup\gamma_{t}$ into finitely many pieces so that $\operatorname{Int}(\gamma_{j})$ contains no vertices in $\{x,x_{1},...,x_{s}\}$ and $\partial\gamma_{j}\in\{a,x_{1},...,x_{s}\}$.
@@
Suppose i=l+1,,si=l+1,...,s. By Lemma 3.15, we have that $\mathring{d}^{\prime}_{i}\geq d^{\prime}=d$. Since $e^{\prime}_{i}\geq\mathring{e}^{\prime}=\mathring{e}$, and $d+\mathring{e}\geq 1$, we conclude that $\mathring{d}^{\prime}_{i}+e^{\prime}_{i}\geq 1$.
Otherwise, let $x_{i},b]_{k}$ be the geodesic in $\mathcal{G}_{k}$ connecting $x_{i}$ and $b$. If $x\in[x_{i},b]_{k}$, then $\mathring{e}_{i}^{\prime}\geq e$. Since $d_{i}^{\prime}\geq\mathring{d}$, we conclude that $\mathring{d}^{\prime}_{i}+e^{\prime}_{i}\geq 1$. Therefore, we may assume that $x\notin[x_{i},b]_{k}$. Then we can break $x_{i},b]_{k}=\gamma_{1}\cup\gamma_{2}\cup...\cup\gamma_{t}$ into finitely many pieces so that $\operatorname{Int}(\gamma_{j})$ contains no vertices in $\{x,x_{1},...,x_{s}\}$ and $\partial\gamma_{j}\in\{b,x_{1},...,x_{s}\}$. We assume that $b\in\partial\gamma_{t}$.
@@
Otherwise, either $j_{n}<i\leq j_{n+1}$ for some n=0,,q1n=0,...,q-1, or $j_{n}\leq i<j_{n+1}$ for some n=p,,1n=p,...,-1. Without loss of generality, we assume that we are in the first case. Let $x_{j_{n}},a]_{k-1}$ and $x_{j_{n+1}},a]_{k-1}$ be the geodesics in $\mathcal{G}_{k-1}$ connecting $x_{j_{n}}$ and $x_{j_{n+1}}$ to $a$ respectively. Note that $\mathring{d}_{j_{n}},\mathring{d}_{j_{n+1}}\leq\mathring{d}_{l}<d$ by Equation (3.13). Thus by Lemma 3.10, $x\notin[x_{j_{n}},a]_{k-1}\cup[x_{j_{n+1}},a]_{k-1}$. Then $\gamma$ must intersect either $x_{j_{n}},a]_{k-1}$ or $x_{j_{n+1}},a]_{k-1}$. Since $\mu_{k-1}$ is $\Gamma_{a,b,k-1}$-admissible, either $\mathring{d}_{j_{n}}+e_{i}^{\prime}\geq 1$ or $\mathring{d}_{j_{n+1}}+e_{i}^{\prime}\geq 1$. Thus,